OpenStudy (gabylovesyou):

Given the equation P2 = A3, what is the orbital period, in years, for the planet Jupiter? (Jupiter is located 5.2 AU from the sun.) 5 years 11 years 110 years 500 years

Miracrown (miracrown):

are those exponents, like P^2 = A^3 maybe?

OpenStudy (perl):

what does A stand for?

Miracrown (miracrown):

and I assume P is the orbital period, but what is A? @Gabylovesyou

Miracrown (miracrown):

I assume we plug in the 5.2 for the A and solve for P but that's just a guess, since I"m not exactly sure what A stands for here.

OpenStudy (turingtest):

i think it is the area swept out over the plane of the ecliptic

Miracrown (miracrown):

ah, that makes sense

OpenStudy (turingtest):

http://www.windows2universe.org/the_universe/uts/kepler3.html good ol' kepler

Miracrown (miracrown):

are we assuming a circular orbit? probably so, since they only give one distance.

OpenStudy (the_fizicx99):

Kepler's 3rd law, I think A is AU, p^2 = 5.2^3 p^2 = 140.608 p = 11.857, I'm not sure if I should round this off.

Miracrown (miracrown):

so, yeah, a = semimajor axis and that's the same as what they give us, the 5.2 so, basically, you just need to cube 5.2 and then take the square root of that answer...

OpenStudy (perl):

, you are using A as radius then

Miracrown (miracrown):

 if I type that in to a calculator, I get something close to one of the choices, but not exactly... I guess we can pick the closest one...